There are the scores of 20 students in a 20-item test in Mathematics in the Modern World: 1, 7, 9, 4, 12, 15, 18, 20, 15, 14, 13, 15, 20, 20, 3, 5, 16, 18, 15 and 13. Find the following:


Answers:

1. Quartile3 =



2. Decile4 =



3. Percentile75 =


1
Expert's answer
2022-01-10T08:01:57-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 x_i & f_i & cf \\
 \hline
 1 & 1 & 1 \\ \hdashline
 3 & 1 & 2 \\ \hdashline
 4 & 1 & 3 \\ \hdashline
 5 & 1 & 4 \\ \hdashline
 7 & 1 & 5 \\ \hdashline
 9 & 1 & 6 \\ \hdashline
 12 & 1 & 7 \\ \hdashline
 13 & 2 & 9 \\ \hdashline
 14 & 1 & 10 \\ \hdashline
 15 & 4 & 14 \\ \hdashline
 16 & 1 & 15 \\ \hdashline
 18 & 2 & 17 \\ \hdashline
 20 & 3 & 20 \\ \hdashline
 & n=20 & \\ 
\end{array}"


1.

"Q_3=(\dfrac{3(n+1) }{4})^{th}" value of the observation

"=(\dfrac{3(20+1) }{4})^{th}" value of the observation

"=(15.75)^{th}" value of the observation


"Q_3=18"


2.

"D_4=(\dfrac{4(n+1) }{10})^{th}" value of the observation

"=(\dfrac{4(20+1) }{10})^{th}" value of the observation

"=(8.4)^{th}" value of the observation


"D_4=13"


3.

"P_{75}=(\dfrac{75(n+1) }{100})^{th}" value of the observation

"=(\dfrac{75(20+1) }{100})^{th}" value of the observation

"=(15.75)^{th}" value of the observation


"P_{75}=18"